3 years ago Lily borrowed 10 000 from faith on condition that she should pay her back two years from now. she also owes faith 6 000 payable five years from now. the applicable interest rate for both transactions is 13.75% per year compounded every six months .after considering her pay back schedule,Lily ask faith if she can buy her 9 000 now and the rest in four years time .she agrees on condition that the new agreement will run from now and that an interest rate of 16.28% per year,compounded monthly, will be applicable from now The amount that Lily will have to pay faith four years from now is


1
Expert's answer
2022-03-09T14:15:42-0500

The amount Lily would have to repay Faith four years from now is given as:

"FV=PV (1+\frac{r}{m})^{mn}"

Where;

PV= (10000+6000)-9000= 7000

r= 16.28% or 0.1628

m= 12

n= 4

"\therefore FV=7000(1+\frac{0.1628}{12})^{12×4}"

"FV=13366.24"

The amount Lily would pay Faith four years from now is 13366.24